OpenStudy (anonymous):

There is no solution for the absolute value inequality [5 x + 2] < - 8 There is a little line under the <

OpenStudy (anonymous):

True or false ?

jhonyy9 (jhonyy9):

I 5x+2 I <= -8 how you solve same inequalitie ?

OpenStudy (anonymous):

@jhonyy9 im not sure what do you mean by solve same inequality ?

OpenStudy (welshfella):

absolute value is always positive

jhonyy9 (jhonyy9):

yes @welshfella but there are 2 ways to solve same inequalite - yes ?

OpenStudy (anonymous):

So its false then ?

jhonyy9 (jhonyy9):

or depend how what way you select to solve it @CC9730 you dont know it again - you need to solve it and will see than that is false or true

jhonyy9 (jhonyy9):

5x+2 <= 8 like a first step subtarct from both sides 8 so what will get ?

jhonyy9 (jhonyy9):

subtract - sorry

jhonyy9 (jhonyy9):

5x+2-8<=8-8

OpenStudy (anonymous):

Am I subtracting 2 or 8 ?

jhonyy9 (jhonyy9):

8 for make the rightb side equal zero

jhonyy9 (jhonyy9):

but you can subtracting 2 too from both sides - yes is right too

jhonyy9 (jhonyy9):

so what will get ?

jhonyy9 (jhonyy9):

5x<=8-2

jhonyy9 (jhonyy9):

5x<=6 divide both sides by 5 x<=6/5 this will be solution of inequalitie

OpenStudy (anonymous):

Sorry im really lost but after dividing well be left with 6/5 ?

jhonyy9 (jhonyy9):

yes so right